OpenStudy (anonymous):

Why does -1/4=-1/x^2 come out as x^2=4? What steps conclude this giving x=2,-2?

OpenStudy (anonymous):

\[-1/4=-1/x^2 \] \[-x^2 = -4\] \[x^2 = 4\] \[x = \sqrt{4}\] \[x = - 2 or x=\]